The following reaction is performed at a certain temperature and these data are acquired. Please use the reaction and data to answer the NExT 2 QUESTONs NO2 (g) H2O2 (aq) No3 (g) H20 (g); AH --16.15 KJ/mol. Kc 144 13. If a scientist observes this reaction at the same temperature starting with 4 M of NO2 and 4 M of TH202, what will be the final concentration of NO3? A. 0.7 M B. 1.8 M C. 3.0 M D. 3.7 M E. 7.4 M 14. Consider the following changes made to the reaction flask in the previous question. Which direction will each of the following changes make the reaction shift? i. increasing temperature ii. increasing the concentration of NO2 iii. decreasing the volume of the container Shift due to i (temp) Shift due to ii (NO2) Shift due to iii (pressure) Right Left Left Left Right Left Left Right Right Right Left Right Right Right
Answer #1

Ans. (D)   

NO2(g) + H2O2(l) <===> NO3(g) + H2O(g)

Con: 4 (1-x) 4x 4x

Kc = [NO3][H2O]/[NO2]

144 = 16x2 /4(1-x)

4x2 +144x -14=0

Using quadratic equation solving

x= 0.9737 M

[NO3] = 4x = 0.9737×4 = 3.8

So answer is D

14)

i) Ans. (A)

Delta H is negative means reaction is exothermic.So increase in temperature shift the reaction to reactant side to nullify the effect of temperature. (Le-Chatlier principle )

ii) B.

increase in [NO2] shift the reaction to product side to nullify the effect of [NO2]. (Le-Chatlier principle )

iii)B (left).

Decrease in volume means increase in no.of moles / unit volume, i.e) pressure increases and reaction proceeds towards backward direction to decrease in no.of moles ,that is towards left.